OpenStudy (anonymous):

Is 3a +1=-3.6(a-1) no solution?

OpenStudy (anonymous):

it looks like there's a solution

OpenStudy (anonymous):

There is?

OpenStudy (anonymous):

3a +1=-3.6(a-1) first distribute 3a +1=-3.6*a-(-3.6)1)

OpenStudy (anonymous):

My teacher said there is no solution for some reason.

OpenStudy (anonymous):

Teach got it wrong this time...:-)

OpenStudy (anonymous):

Yeah it is no Solution

OpenStudy (anonymous):

3a +1=-3.6(a-1) 3a +1=-3.6a+3.6) 6.6a=2.6 a=26/66 a=13/33 how is it no solution? @samdhimal

OpenStudy (anonymous):

dude @completeidiot The value of a is -1 are you blind

OpenStudy (anonymous):

i dont see an a = -1

OpenStudy (anonymous):

i mean a-1 that means a=1

OpenStudy (anonymous):

use that and See what happen

OpenStudy (anonymous):

then it wouldnt make sense, why would you have an equation within an expression i see a minus 1 not a is equal to 1 - and = are 2 different things

OpenStudy (anonymous):

LOl you have to think about it

OpenStudy (anonymous):

a=1 only if a-1=0 which is not true